# Practical hybrid decoding scheme for parity-encoded spin systems

### v3

@@ III.1 Probabilistic decoding
The second term of Eq.(13) enforces the parity constraint that $\boldsymbol{z}$ should obey Eq.(5) to be a valid code-state. In contrast, the first term captures the correlation between the observation $\boldsymbol{y}$ and the code-state $\boldsymbol{z}$. The latter formulation is equivalent to the former one as long as $\boldsymbol{z}$ is a valid code-state associated with a source-state $\boldsymbol{Z}$, i.e., $s_{i}(\boldsymbol{z})=1$ for any $i$. We note that $H^{code}\left(\boldsymbol{z}\right)$ can be considered a Hamiltonian of an enlarged spin system. In this formulation, the word MAP decoding corresponds to finding the most probable code-state $\boldsymbol{z}$, namely,
$\boldsymbol{z}=\underset{\boldsymbol{x}\in C}{\arg\max}P(\boldsymbol{x}|\boldsymbol{y})=\underset{\boldsymbol{x}\in C}{\arg\min}H^{code}(\boldsymbol{x}),$ (15)
where $C$ denotes the set of all code-states. Such decoded results can be obtained, for example, by SA and/or QA.
Now, let us recall the SLHZ model. In this case, $s_{i}(\boldsymbol{z})$ is given by the weight-4 syndrome in the spin representation, which is written as a product of four elements in $\boldsymbol{z}$, i.e., $s_{klmn}^{(4)}(\boldsymbol{z})=z_{km}z_{lm}z_{ln}z_{kn}$ [10]. Then, we see that $H^{code}\left(\boldsymbol{z}\right)$ agrees with the Hamiltonian of the SLHZ model. Therefore, the word MAP decoding of the LDPC codes is equivalent to finding the ground state of the SLHZ model. Furthermore, since the two formulations, i.e., those based on $H^{code}\left(\boldsymbol{z}\right)$ (Eq.(13)) and those based on $H^{source}\left(\boldsymbol{Z}\right)$ (Eq.(8)) are mathematically equivalent, it follows that finding the ground code-state of the SLHZ model is equivalent to finding the ground state of the spin glass.
Symbol MAP decoding
The above discussion implies that we can solve the COPs through decoding the associated classical LDPC codes. Meanwhile, the word MAP decoding is not the only strategy for decoding classical LDPC codes. There is another decoding strategy based on different principles. Instead of considering the most probable word, it is also allowed to be interested only in the most probable symbol, i.e., the most probable value $z_{i}$ of the $i$th spin, ignoring the values of the other spin variables. In this strategy, we consider marginals,
$P(x_{i}|\boldsymbol{y})$ $:=$ $\sum_{x_{1}=-1}^{+1}\cdots\sum_{x_{i-1}=-1}^{+1}\sum_{x_{i+1}=-1}^{+1}\cdots\sum_{x_{N_{v}}=-1}^{+1}P(\boldsymbol{x}|\boldsymbol{y})$ (16)
$=$ $\sum_{x_{k}(k\neq i)}P(\boldsymbol{x}|\boldsymbol{y}).$
Decoding corresponds to finding the value of the spin variable that maximizes the marginals $P\left(x_{i}|\boldsymbol{y}\right)$, i.e.,
$z_{i}$ $=$ $\underset{x_{i}\in\{\pm 1\}}{\arg\max}P(x_{i}|\boldsymbol{y})$ (17)
$=$ $\mathrm{sign}[P(x_{i}=+1|\boldsymbol{y})-P(x_{i}=-1|\boldsymbol{y})]$
$=$ $\mathrm{sign}[L_{i}(\boldsymbol{y})],$
@@
$L_{i}(\boldsymbol{y})=\log\frac{P(x_{i}=+1|\boldsymbol{y})}{P(x_{i}=-1|\boldsymbol{y})}$ (18)
is called the a posteriori LLR for the $i$th spin. Since the absolute value $|L_{i}(\boldsymbol{y})|$ gives the error probability of this decision as
@@
it represents a metric to measure the reliability or uncertainty of a decision in Eq.(17), which indicates how likely $x_{i}$ it is to be +1+1 or 1-1. A value close to zero indicates a bad inference, while a larger values indicate a better inference. [27].
A variety of algorithms can achieve symbol MAP decoding. For example, the belief propagation (BP) algorithm is widely regarded as the best-known algorithm for LDPC codes [21]. The BP algorithm is an iterative, deterministic algorithm in which the a posteriori LLRs, initially given by the observation $\boldsymbol{y}$, gradually increase in absolute value as messages are iteratively exchanged between the VNs and CNs.
In the following section, we study an alternative, simple, and straightforward algorithm: the bit-flipping (BF) algorithm [22, 23]. This algorithm is a variety of the symbol MAP decoding strategies. It uses the syndrome vector $\boldsymbol{s}(\boldsymbol{x})=(s_{1}(\boldsymbol{x}),\ldots,s_{N_{c}}(\boldsymbol{x}))$ as a key to decide the most probable value $z_{i}$ of the $i$th spin, namely,
$z_{i}=\underset{x_{i}\in\{\pm 1\}}{\arg\max}P(x_{i}|\boldsymbol{y})=\mathrm{sign}[\tilde{L}_{i}(\boldsymbol{x})],$ (20)
where $\boldsymbol{x}=\mathrm{sign}[\boldsymbol{y}]\in\{\pm 1\}^{N_{v}}$ is the hard decision of the observation $\boldsymbol{y}$ and
$\tilde{L}_{i}(\boldsymbol{x})=\log\frac{P(z_{i}=+1|\boldsymbol{B}_{i}(\boldsymbol{x}))}{P(z_{i}=-1|\boldsymbol{B}_{i}(\boldsymbol{x}))}$ (21)
is the associated a posteriori LLR for the $i$th spin, where $\boldsymbol{B}_{i}(\boldsymbol{x})$ is the estimator of $z_{i}$ which depends on $\boldsymbol{s}(\boldsymbol{x})$, as shown later. The algorithm starts with an initial decision $\boldsymbol{x}^{(0)}=\mathrm{sign}[\boldsymbol{y}]$, and iteratively updates the decision $x_{i}^{(m)}=\mathrm{sign}[\tilde{L}_{i}(\boldsymbol{x}^{(m-1)})]$ by a spin-flip operation, which depends on the decision $\boldsymbol{x}^{(m-1)}=(x_{1}^{(m-1)},\ldots,x_{N_{v}}^{(m-1)})$ in the previous round. If the decision $\boldsymbol{x}$ converges after several rounds of iteration, we consider it the most probable code-state $\boldsymbol{z}$.
In the BF algorithm, the reliability of the decision $|\tilde{L}_{i}(\boldsymbol{x})|$ gradually increases as the spin-flipping is repeated. It is widely recognized that one of the key advantages of the BF algorithm over the BP algorithm is its simplicity and lower computational complexity. The BF algorithm is particularly useful in environments with limited computational resources because it requires fewer operations to decode the received message. This fact makes it a more efficient option when processing power is a concern. In contrast, although it provides better error-correcting performance in terms of bit error rate (BER), the BP algorithm typically involves more complex calculations and higher computational requirements. Therefore, the BF decoding is often favored in communication applications where simplicity and low complexity are prioritized.
@@ III.2 Readout decoding of the SLHZ model
Bit flipping decoding algorithm
@@
Now, let us analyze readout decoding of the SLHZ model from the perspective of the LDPC codes. In the following analysis, we assume that the readout $\boldsymbol{r}=(r_{12},\ldots,r_{K-1\,K})\in\left\{\pm 1\right\}^{\tbinom{K}{2}}$ is available for the $\tbinom{K}{2}$ physical spins in the SLHZ model in accordance with the Ising model. We disregard the information about the observations $\boldsymbol{y}$ of the physical spins, even if it is available in the actual implementation. We will revisit this point later.
@@
We first consider the simplest case, i.e., correcting errors caused by i.i.d. noise. This model is appropriate for an SLHZ-based QA device when measurement errors dominate other readout errors. We propose a simple decoding algorithm with very low decoding complexity, which is associated with Gallager’s BF decoding algorithm in the spin representation [22, 23]. It is based on Massey’s APP (a posteriori probability) algorithm [27]. Our algorithm only uses information about the syndrome vector $\boldsymbol{s}(\boldsymbol{r})$ to estimate errors $\boldsymbol{e}$ in the current decision $\boldsymbol{r}$ for the readout. Let us introduce the error pattern by $\boldsymbol{e}=\boldsymbol{r}\circ\boldsymbol{z}=(e_{12},\ldots,e_{K-1\,K})\in\left\{\pm 1\right\}^{\tbinom{K}{2}}$, where $\circ$ denotes componentwise multiplication. The syndrome vector depends only on $\boldsymbol{e}$ because
$\boldsymbol{s}(\boldsymbol{z})=\left(\prod_{\left\{j:H_{1j}=1\right\}}z_{j},\ldots,\prod_{\left\{j:H_{N_{c}j}=1\right\}}z_{j}\right)=(+1,\ldots,+1)$ (22)
@@
holds for any code-state $\boldsymbol{z}$. Therefore, if we note that $z_{ij}\in\left\{\pm 1\right\}$ and $e_{ij}\in\left\{\pm 1\right\}$,
$\boldsymbol{s}(\boldsymbol{r})=\boldsymbol{s}(\boldsymbol{z}\circ\boldsymbol{e})=\boldsymbol{s}(\boldsymbol{z})\circ\boldsymbol{s}(\boldsymbol{e})=\boldsymbol{s}(\boldsymbol{e})$ (23)
@@
holds. It is important to note that although we never know $\boldsymbol{e}$, we can know $\boldsymbol{s}(\boldsymbol{e})$ from the knowledge of the decision $\boldsymbol{r}$ and utilize it to estimate errors $\boldsymbol{e}$ to be eliminated.
@@
Estimating the correct value $z_{ij}$ is equivalent to estimating error $e_{ij}$, as $z_{ij}=r_{ij}e_{ij}$ holds. The algorithm uses the weight-3 syndrome $s_{ijk}^{(3)}\left(\boldsymbol{r}\right)=r_{ij}r_{jk}r_{ik}=e_{ij}e_{jk}e_{ik}$ rather than the weight-4 syndrome to decode the current decision $\boldsymbol{r}$ because it has the advantage of treating all variables (VN and CN) symmetrically [20]. The trade-off is that more constraints may increase the decoding complexity. However, as will be discussed later, this is not a fatal problem, as we can leverage symmetry to reduce computational costs. An optimal estimate $e_{ij}^{*}$ can be obtained from $N_{v}=\tbinom{K}{2}$ syndromes $s_{ijk}^{(3)}\left(\boldsymbol{r}\right)$ for $k\neq i,j$. They are considered K2K-2 parity checks orthogonal on $e_{ij}$ [27] and their values suggest whether the parity-check equation is satisfied (being +1+1) or violated (being 1-1). Based on the majority voting, we decide an error variable $e_{ij}^{*}$ whether the spin $ij$ should be flipped ($e_{ij}^{*}=-1$) or not ($e_{ij}^{*}=+1$). The optimal estimate $e_{ij}^{*}$ can be obtained according to the APP decoding by [27]
$e_{ij}^{*}=\mathrm{sign}[\Delta_{ij}(\boldsymbol{r})],$ (24)
@@
$\Delta_{ij}(\boldsymbol{r})=w_{0}+\sum_{k\neq i,j}^{K}w_{k}s_{ijk}^{(3)}(\boldsymbol{r})$ (25)
@@
$r_{ij}^{*}=r_{ij}e_{ij}^{*}=\mathrm{sign}[L_{ij}^{*}(\boldsymbol{r})],$ (30)
@@
$L_{ij}^{*}(\boldsymbol{r})$ $=$ $r_{ij}\Delta_{ij}(\boldsymbol{r})=\log\frac{P(r_{ij}^{*}=+1|\boldsymbol{B}_{ij}(\boldsymbol{r})}{P(r_{ij}^{*}=-1|\boldsymbol{B}_{ij}(\boldsymbol{r}))},$ (31)
$\boldsymbol{B}_{ij}(\boldsymbol{r})$ $=$ $(B_{ij0}(\boldsymbol{r}),B_{ij1}(\boldsymbol{r}),\ldots,B_{ijK}(\boldsymbol{r}))$ (32)
$=$ $r_{ij}(1,s_{ij1}^{(3)}(\boldsymbol{r}),\ldots,s_{ijK}^{(3)}(\boldsymbol{r})),$
@@
$\gamma_{ij}$ $=$ $P(r_{ij}\neq r_{ij}^{*})=P(B_{ij0}(\boldsymbol{r})\neq r_{ij}^{*}),$ (33)
$p_{k}$ $=$ $P(r_{ij}s_{ijk}^{(3)}(\boldsymbol{r})\neq r_{ij}^{*})=P(B_{ijk}(\boldsymbol{r})\neq r_{ij}^{*}).$ (34)
holds [27, 30]. The function $L_{ij}^{*}(\boldsymbol{r})$ can be considered as the LLR for the decision $r_{ij}^{*}$ based on the estimators $\boldsymbol{B}_{ij}(\boldsymbol{r})$, which depends on the weight-3 syndrome vector $\boldsymbol{s}^{(3)}(\boldsymbol{r})$.
Figure 4: Binary symmetric channel associated with the parameters $\gamma_{ij}$ and $p_{k}$.
In Eq.(24), $e_{ij}^{*}=-1$ indicates that we should invert the sign of $r_{ij}\rightarrow r_{ij}^{*}=r_{ij}e_{ij}^{*}=-r_{ij}$. Parameters $\gamma_{ij}$ and $p_{k}$ are crosstalk parameters that characterize the binary symmetric channels shown in Fig.4. We should note that the probability $p_{k}$ is given by the probability of an odd number of 1-1’s among the errors exclusive of $e_{ij}$ that are checked by $s_{ijk}^{(3)}\left(\boldsymbol{r}\right)=e_{ij}e_{jk}e_{ik}$ so that it is given by [27]
@@
If we assume i.i.d. noise, $\gamma_{ij}=\gamma_{0}$ holds for every $\left\{i,j\right\}$ where $0<\gamma_{0}<\tfrac{1}{2}$ , we can confirm that $w_{0}$ can be approximated $w_{0}=w_{k}>0$ for any $k$. Then, we obtain the following simple algorithm: we calculate the best estimate $e_{ij}^{*}$ by
$e_{ij}^{*}=\mathrm{sign}\left(1+\sum_{k\neq i,j}^{K}s_{ijk}^{(3)}(\boldsymbol{r})\right).$ (36)
Figure 5: Bipartite graph for $K=5$ logical spins. Solid blue lines show an example of the shortest loop of edges connecting VNs and CNs, which has a length of 6.
This equation means that if the majority vote of the estimators given by $\boldsymbol{A}_{ij}(\boldsymbol{r})=\left(1,s_{ij1}^{(3)}(\boldsymbol{x}),\ldots,s_{ijK}^{(3)}(\boldsymbol{x})\right)$ is negative, $r_{ij}$ should be flipped to increase the value of $\sum_{k\neq i,j}^{K}s_{ijk}^{(3)}(\boldsymbol{r})$. Thus, Eq.(36) is reduced to the following best estimate $r_{ij}^{*}\in\left\{\pm 1\right\}$:
@@
Eq.(37) can be interpreted as Gallager’s BF decoding [22, 23]. Note that the best estimate $r_{ij}^{*}$ is represented in terms of the values of the current variables in VNs. Fig.5 shows the relevant bipartite graph for $K=5$. Note that the graph associated with the SLHZ system is more loopy, with a minimum length of 4, whereas this graph is less loopy, with a minimum length of 6. The initial decision $\boldsymbol{r}$ on the VNs is broadcast to its adjacent CNs connected by edges. Each CN then reports the syndrome value to its adjacent VNs connected by the edge. All VNs update their values simultaneously according to Eq.(36), representing a majority vote of 1 and the values of adjacent CNs.
@@
Next, let us discuss how to apply the BF decoding to the SLHZ model. We introduce a matrix representation of the state of the SLHZ model. Consider the $K\times K$ symmetrized matrix $\hat{\boldsymbol{z}}$ whose elements are given by spin variable $z_{ij}=z_{ji}\in\left\{\pm 1\right\}$ and have unit diagonal elements, namely,
$\hat{\boldsymbol{z}}$ $=$ $\left[\begin{array}[]{cccccc}1&z_{12}&z_{13}&\cdots&z_{1\,K-1}&z_{1\,K}\\ z_{21}&1&z_{23}&\cdots&z_{2\,K-1}&z_{2\,K}\\ z_{31}&z_{32}&1&\cdots&z_{3\,K-1}&z_{3\,K}\\ \vdots&\vdots&\vdots&\ddots&\cdots&\vdots\\ z_{K-1\,1}&z_{K-1\,2}&z_{K-1\,3}&\cdots&1&z_{K-1\,K}\\ z_{K\,1}&z_{K\,2}&z_{K\,3}&\cdots&z_{K\,K-1}&1\end{array}\right].$ (44)
We use similar notations $\hat{\boldsymbol{r}}$ and $\hat{\boldsymbol{e}}$ for the symmetrized matrices associated with the current readout $\boldsymbol{r}$ and error pattern $\boldsymbol{e}$. They satisfy $\hat{\boldsymbol{r}}=\hat{\boldsymbol{z}}\circ\hat{\boldsymbol{e}}$. Hereafter, we will refer to $\hat{\boldsymbol{e}}$ as the error matrix. Then, Eq.(37) can be conveniently written as
$\hat{\boldsymbol{r}}^{*}=\mathcal{F}(\hat{\boldsymbol{r}})=\mathrm{sign}\left[\hat{\boldsymbol{r}}(\hat{\boldsymbol{r}}-\boldsymbol{I}_{K\times K})\right],$ (46)
where we assume that the sign function is componentwise. This operation updates the $\tbinom{K}{2}$ elements in $\hat{\boldsymbol{r}}$ simultaneously. Therefore, Eq.(46) represents a parallel BF algorithm. We further consider the iterative operation of $\mathcal{F}$, i.e.,

### v4

@@ III.1 Probabilistic decoding
The second term of Eq.(13) enforces the parity constraint that z should obey Eq.(5) to be a valid code state. In contrast, the first term captures the correlation between the observation y and the code state z. The latter formulation is equivalent to the former one as long as z is a valid code state associated with a source state Z, i.e., $s_{i}(\textbf{z})=1$ for any $i$. We note that $H^{code}\left(\textbf{z}\right)$ can be considered a Hamiltonian of an enlarged spin system. In this formulation, the word-MAP decoding corresponds to finding the most probable code state z, namely,
$\textbf{z}=\underset{\textbf{x}\in C}{\arg\max}P(\textbf{x}|\textbf{y})=\underset{\textbf{x}\in C}{\arg\min}H^{code}(\textbf{x}),$ (15)
where $C$ denotes the set of all code states. Such decoded results can be obtained by, e.g., SA and/or QA.
Now, let us recall the SLHZ model. In this case, $s_{i}(\textbf{z})$ is given by the weight-4 syndrome in the spin representation, which is written as a product of four elements in $\boldsymbol{z}$, i.e., $s_{klmn}^{(4)}(\textbf{z})=z_{km}z_{lm}z_{ln}z_{kn}$ [10]. Then, we see that $H^{code}\left(\textbf{z}\right)$ agrees with the Hamiltonian of the SLHZ model. Therefore, the word-MAP decoding of the LDPC codes is equivalent to finding the ground state of the SLHZ model. Furthermore, since the two formulations, i.e., those based on $H^{code}\left(\textbf{z}\right)$ (Eq.(13)) and those based on $H^{source}\left(\textbf{Z}\right)$ (Eq.(8)) are mathematically equivalent, it follows that finding the ground code state of the SLHZ model is equivalent to finding the ground state of the spin glass.
Symbol-MAP decoding
The above discussion implies that we can solve the COPs through decoding the associated classical LDPC codes. Meanwhile, the word-MAP decoding is not the only strategy for decoding classical LDPC codes. There is another decoding strategy based on different principles. Instead of considering the most probable word, it is also allowed to be interested only in the most probable symbol, i.e., the most probable value $z_{i}$ of the $i$th spin, ignoring the values of the other spin variables. In this strategy, we consider marginals,
$P(x_{i}|\textbf{y})$ $:=$ $\sum_{x_{1}=-1}^{+1}\cdots\sum_{x_{i-1}=-1}^{+1}\sum_{x_{i+1}=-1}^{+1}\cdots\sum_{x_{N_{v}}=-1}^{+1}P(\textbf{x}|\textbf{y})$ (16)
$=$ $\sum_{x_{k}(k\neq i)}P(\textbf{x}|\textbf{y}).$
Decoding corresponds to finding the value of the spin variable that maximizes the marginals $P\left(x_{i}|\textbf{y}\right)$, i.e.,
$z_{i}$ $=$ $\underset{x_{i}\in\{\pm 1\}}{\arg\max}P(x_{i}|\textbf{y})$ (17)
$=$ $\mathrm{sign}[P(x_{i}=+1|\textbf{y})-P(x_{i}=-1|\textbf{y})]$
$=$ $\mathrm{sign}[L_{i}(\textbf{y})],$
@@
$L_{i}(\textbf{y})=\log\frac{P(x_{i}=+1|\textbf{y})}{P(x_{i}=-1|\textbf{y})}$ (18)
is called the a posteriori LLR for the $i$th spin. Since the absolute value $|L_{i}(\textbf{y})|$ gives the error probability of this decision as
@@
it represents a metric to measure the reliability or uncertainty of a decision in Eq.(17), which indicates how likely $x_{i}$ it is to be +1+1 or 1-1. A value close to zero indicates a bad inference, while a larger value indicates a better inference. [26].
A variety of algorithms can achieve symbol-MAP decoding. For example, the BP algorithm is widely regarded as the best-known algorithm for LDPC codes [21]. The BP algorithm is an iterative, deterministic algorithm in which the a posteriori LLRs, initially given by the observation y, gradually increase in absolute value as messages are iteratively exchanged between the VNs and CNs.
In the following section, we study an alternative, simple, and straightforward algorithm: the BF algorithm [22, 23]. This algorithm is a variety of the symbol-MAP decoding strategies. It uses the syndrome vector $\textbf{s}(\textbf{x})=(s_{1}(\textbf{x}),\ldots,s_{N_{c}}(\textbf{x}))$ as a key to decide the most probable value $z_{i}$ of the $i$th spin, namely,
$z_{i}=\underset{x_{i}\in\{\pm 1\}}{\arg\max}P(x_{i}|\textbf{y})=\mathrm{sign}[\tilde{L}_{i}(\textbf{x})],$ (20)
where $\textbf{x}=\mathrm{sign}[\textbf{y}]\in\{\pm 1\}^{N_{v}}$ is the hard decision of the observation y and
$\tilde{L}_{i}(\textbf{x})=\log\frac{P(z_{i}=+1|\boldsymbol{B}_{i}(\textbf{x}))}{P(z_{i}=-1|\boldsymbol{B}_{i}(\textbf{x}))}$ (21)
is the associated a posteriori LLR for the $i$th spin, where $\textbf{B}_{i}(\textbf{x})$ is the estimator of $z_{i}$ which depends on $\textbf{s}(\textbf{x})$, as shown later. The algorithm starts with an initial decision $\textbf{x}^{(0)}=\mathrm{sign}[\textbf{y}]$, and iteratively updates the decision $x_{i}^{(m)}=\mathrm{sign}[\tilde{L}_{i}(\textbf{x}^{(m-1)})]$ by a spin-flip operation, which depends on the decision $\textbf{x}^{(m-1)}=(x_{1}^{(m-1)},\ldots,x_{N_{v}}^{(m-1)})$ in the previous round. If the decision x converges after several rounds of iteration, we consider it the most probable code state z.
In the BF algorithm, the reliability of the decision $|\tilde{L}_{i}(\textbf{x})|$ gradually increases as the spin flipping is repeated. It is widely recognized that one of the key advantages of the BF algorithm over the BP algorithm is its simplicity and lower computational complexity. The BF algorithm is particularly useful in environments with limited computational resources because it requires fewer operations to decode the received message. This fact makes it a more efficient option when processing power is a concern. In contrast, although it provides better error-correcting performance in terms of the bit error rate (BER), the BP algorithm typically involves more complex calculations and higher computational requirements. Therefore, the BF decoding is often favored in communication applications where simplicity and low complexity are prioritized.
@@ III.2 Readout decoding of the SLHZ model
Bit-flipping decoding algorithm
@@
Now, let us analyze readout decoding of the SLHZ model from the perspective of the LDPC codes. In the following analysis, we assume that the readout $\textbf{r}=(r_{12},\ldots,r_{K-1\,K})\in\left\{\pm 1\right\}^{\tbinom{K}{2}}$ is available for the $\tbinom{K}{2}$ physical spins in the SLHZ model in accordance with the Ising model. We disregard the information about the observations y of the physical spins, even if it is available in the actual implementation. We will revisit this point later.
@@
We first consider the simplest case, i.e., correcting errors caused by IID noise. This model is appropriate for an SLHZ-based QA device when measurement errors dominate other readout errors. We propose a simple decoding algorithm with very low decoding complexity, which is associated with Gallager’s BF decoding algorithm in the spin representation [22, 23]. It is based on Massey’s a posteriori probability (APP) algorithm [26]. Our algorithm only uses information about the syndrome vector $\textbf{s}(\textbf{r})$ to estimate errors e in the current decision r for the readout. Let us introduce the error pattern by $\textbf{e}=\textbf{r}\circ\textbf{z}=(e_{12},\ldots,e_{K-1\,K})\in\left\{\pm 1\right\}^{\tbinom{K}{2}}$, where $\circ$ denotes component-wise multiplication. The syndrome vector depends only on e because
$\textbf{s}(\textbf{z})=\left(\prod_{\left\{j:H_{1j}=1\right\}}z_{j},\ldots,\prod_{\left\{j:H_{N_{c}j}=1\right\}}z_{j}\right)=(+1,\ldots,+1)$ (22)
@@
holds for any code state z. Therefore, if we note that $z_{ij}\in\left\{\pm 1\right\}$ and $e_{ij}\in\left\{\pm 1\right\}$,
$\textbf{s}(\textbf{r})=\textbf{s}(\textbf{z}\circ\textbf{e})=\textbf{s}(\textbf{z})\circ\textbf{s}(\textbf{e})=\textbf{s}(\textbf{e})$ (23)
@@
holds. It is important to note that although we never know e, we can know $\textbf{s}(\textbf{e})$ from the knowledge of the decision r and utilize it to estimate errors e to be eliminated.
@@
Estimating the correct value $z_{ij}$ is equivalent to estimating error $e_{ij}$, as $z_{ij}=r_{ij}e_{ij}$ holds. The algorithm uses the weight-3 syndrome $s_{ijk}^{(3)}\left(\textbf{r}\right)=r_{ij}r_{jk}r_{ik}=e_{ij}e_{jk}e_{ik}$ rather than the weight-4 syndrome to decode the current decision r because it has the advantage of treating all variables (VN and CN) symmetrically [20]. The trade-off is that more constraints may increase the decoding complexity. However, as will be discussed later, this is not a fatal problem, as we can leverage symmetry to reduce computational costs. An optimal estimate $e_{ij}^{*}$ can be obtained from $N_{v}=\tbinom{K}{2}$ syndromes $s_{ijk}^{(3)}\left(\textbf{r}\right)$ for $k\neq i,j$. They are considered K2K-2 parity checks orthogonal on $e_{ij}$ [26] and their values suggest whether the parity-check equation is satisfied (being +1+1) or violated (being 1-1). Based on the majority voting, we decide an error variable $e_{ij}^{*}$ whether the spin $ij$ should be flipped ($e_{ij}^{*}=-1$) or not ($e_{ij}^{*}=+1$). The optimal estimate $e_{ij}^{*}$ can be obtained according to the APP decoding by [26]
$e_{ij}^{*}=\mathrm{sign}[\Delta_{ij}(\textbf{r})],$ (24)
@@
$\Delta_{ij}(\textbf{r})=w_{0}+\sum_{k\neq i,j}^{K}w_{k}s_{ijk}^{(3)}(\textbf{r})$ (25)
@@
$r_{ij}^{*}=r_{ij}e_{ij}^{*}=\mathrm{sign}[L_{ij}^{*}(\textbf{r})],$ (30)
@@
$L_{ij}^{*}(\textbf{r})$ $=$ $r_{ij}\Delta_{ij}(\textbf{r})=\log\frac{P(r_{ij}^{*}=+1|\textbf{B}_{ij}(\textbf{r})}{P(r_{ij}^{*}=-1|\textbf{B}_{ij}(\textbf{r}))},$ (31)
$\textbf{B}_{ij}(\textbf{r})$ $=$ $(B_{ij0}(\textbf{r}),B_{ij1}(\textbf{r}),\ldots,B_{ijK}(\textbf{r}))$ (32)
$=$ $r_{ij}(1,s_{ij1}^{(3)}(\textbf{r}),\ldots,s_{ijK}^{(3)}(\textbf{r})),$
@@
$\gamma_{ij}$ $=$ $P(r_{ij}\neq r_{ij}^{*})=P(B_{ij0}(\textbf{r})\neq r_{ij}^{*}),$ (33)
$p_{k}$ $=$ $P(r_{ij}s_{ijk}^{(3)}(\textbf{r})\neq r_{ij}^{*})=P(B_{ijk}(\textbf{r})\neq r_{ij}^{*}).$ (34)
holds [26, 29]. The function $L_{ij}^{*}(\textbf{r})$ can be considered as the LLR for the decision $r_{ij}^{*}$ based on the estimators $\textbf{B}_{ij}(\textbf{r})$, which depends on the weight-3 syndrome vector $\textbf{s}^{(3)}(\textbf{r})$.
Figure 4: The binary symmetric channel associated with the parameters $\gamma_{ij}$ and $p_{k}$.
In Eq.(24), $e_{ij}^{*}=-1$ indicates that we should invert the sign of $r_{ij}\rightarrow r_{ij}^{*}=r_{ij}e_{ij}^{*}=-r_{ij}$. Parameters $\gamma_{ij}$ and $p_{k}$ are crosstalk parameters that characterize the binary symmetric channels shown in Fig.4. We should note that the probability $p_{k}$ is given by the probability of an odd number of 1-1’s among the errors exclusive of $e_{ij}$ that are checked by $s_{ijk}^{(3)}\left(\textbf{r}\right)=e_{ij}e_{jk}e_{ik}$ so that it is given by [26]
@@
If we assume IID noise, such that $\gamma_{ij}=\gamma_{0}$ holds for every $\{i,j\}$ where $0<\gamma_{0}<\tfrac{1}{2}$ , we can confirm that $w_{0}$ can be approximated $w_{0}=w_{k}>0$ for any $k$. Then, we obtain the following simple algorithm: we calculate the best estimate $e_{ij}^{*}$ by
$e_{ij}^{*}=\mathrm{sign}\left(1+\sum_{k\neq i,j}^{K}s_{ijk}^{(3)}(\textbf{r})\right).$ (36)
Figure 5: The bipartite graph for $K=5$ logical spins. The solid blue lines show an example of the shortest loop of edges connecting VNs and CNs, which has a length of 6.
This equation means that if the majority vote of the estimators given by $\textbf{A}_{ij}(\textbf{r})=\left(1,s_{ij1}^{(3)}(\textbf{x}),\ldots,s_{ijK}^{(3)}(\textbf{x})\right)$ is negative, $r_{ij}$ should be flipped to increase the value of $\sum_{k\neq i,j}^{K}s_{ijk}^{(3)}(\textbf{r})$. Thus, Eq.(36) is reduced to the following best estimate $r_{ij}^{*}\in\left\{\pm 1\right\}$:
@@
Eq.(37) can be interpreted as Gallager’s BF decoding [22, 23]. Note that the best estimate $r_{ij}^{*}$ is represented in terms of the values of the current variables in VNs. In Fig.5, we shows the relevant bipartite graph for $K=5$. Note that the graph associated with the SLHZ system is more loopy, with a minimum length of 4, whereas this graph is less loopy, with a minimum length of 6. The initial decision r on the VNs is broadcast to its adjacent CNs connected by edges. Each CN then reports the syndrome value to its adjacent VNs connected by the edge. All VNs update their values simultaneously according to Eq.(36), representing a majority vote of $1$ and the values of adjacent CNs.
@@
Next, let us discuss how to apply the BF decoding to the SLHZ model. We introduce a matrix representation of the state of the SLHZ model. Consider the $K\times K$ symmetrized matrix $\hat{\textbf{z}}$, the elements of which are given by spin variable $z_{ij}=z_{ji}\in\left\{\pm 1\right\}$ and have unit diagonal elements, namely,
$\hat{\textbf{z}}$ $=$ $\left[\begin{array}[]{cccccc}1&z_{12}&z_{13}&\cdots&z_{1\,K-1}&z_{1\,K}\\ z_{21}&1&z_{23}&\cdots&z_{2\,K-1}&z_{2\,K}\\ z_{31}&z_{32}&1&\cdots&z_{3\,K-1}&z_{3\,K}\\ \vdots&\vdots&\vdots&\ddots&\cdots&\vdots\\ z_{K-1\,1}&z_{K-1\,2}&z_{K-1\,3}&\cdots&1&z_{K-1\,K}\\ z_{K\,1}&z_{K\,2}&z_{K\,3}&\cdots&z_{K\,K-1}&1\end{array}\right].$ (44)
We use similar notations $\hat{\textbf{r}}$ and $\hat{\textbf{e}}$ for the symmetrized matrices associated with the current readout r and error pattern e. They satisfy $\hat{\textbf{r}}=\hat{\textbf{z}}\circ\hat{\textbf{e}}$. Hereafter, we will refer to $\hat{\textbf{e}}$ as the error matrix. Then, Eq.(37) can be conveniently written as
$\hat{\textbf{r}}^{*}=\mathcal{F}(\hat{\textbf{r}})=\mathrm{sign}\left[\hat{\textbf{r}}(\hat{\textbf{r}}-\textbf{I}_{K\times K})\right],$ (46)
where we assume that the sign function is component-wise. This operation updates the $\tbinom{K}{2}$ elements in $\hat{\textbf{r}}$ simultaneously. Therefore, Eq.(46) represents a parallel BF algorithm. We further consider the iterative operation of $\mathcal{F}$, i.e.,